OpenStudy (anonymous):

finding means, variances and correlation coefficient of X and Y of p(x,y)=e^-2/[x!(y-x)!] , y=0,1,2... x=0,1,...y

OpenStudy (anonymous):

i have found the mgf of this joint distribution it is \[e ^{-2}e ^{e _{}^{t}(e ^{s}+1)}\]

OpenStudy (zarkon):

you found the hardest part of this problem...where are you stuck?

OpenStudy (anonymous):

to find the mean take derivative but with respect to what? and set to 0

OpenStudy (zarkon):

\[M_{X,Y}(s,t)=E\left[e^{sX+tY}\right]\] \[\frac{d}{dt}E\left[e^{sX+tY}\right]=E\left[Xe^{sX+tY}\right]\] set s,t =0 \[=E\left[Xe^{0X+0Y}\right]=E[X]\]